# Fully automated workflow for the design of patient-specific orthopaedic implants: application to total knee arthroplasty

### v2

@@ 2.5.1 Femoral implant
The design of the femoral implant starts with the posterior part and in first place the medial pivot sphere (Figure 4a). For each condyle, the radius of the medial-pivot sphere is computed from the antero-posterior size [41]. Both the medial and lateral post-condylar surfaces are initially shaped based on this sphere, after which they are refined to tightly align with the contours of the resected bone. The posterior most proximal part is then cut along another circle arc to obtain the roll back shape that will allow deep flexion. Secondly, the anterior surface is designed to closely adhere to the patient’s premorbid anatomy, aiming to facilitate patellar tracking. The specific characteristics of each patient’s lateral and medial facet height and sulcus angle are individually taken into consideration to prevent the risk of patellar instability [42, 43].
@@ 2.5.2 Tibial implant
The shape of the tibial baseplate is crafted to conform to the contours of the resected bone, ensuring optimal contact with cortical bone for enhanced implant stability and durability (Figure 4). The location of the keel is determined by taking into account the morphology of the diaphysis and the tibial plateau. The length of the stem, as well as the dimensions of the two vanes, are also functions of the patient’s specific characteristics.
@@
Figure 4: Design steps for the personalised prosthesis. a) Femoral implant, from left to right: defining the medial pivot sphere; defining the condyles’ curvature; defining the anterior surface; cutting the intercondylar notch; adding pins and chamfers. b) Tibial implant, from left to right: tibial baseplate contouring; adding locking mechanism for the polyethylene insert; adding the keel; adding chamfers.
@@ 2.5.3 Tibial insert
The polyethylene tibial insert is designed regarding the femoral and tibial implants. The basic shape of the insert is built from the embase shape of the tibial component. The footprint of the medial pivot sphere is excavated into the basic shape. The lateral part is designed almost flat, to allow free antero-posterior translation of the femoral component as suggested by the medial-pivot concept. Finally the most anterior part is cut to avoid conflict with the patella in deep flexion.
@@ 2.5.4 Design evaluation
To evaluate the quality of the designed custom implants, we computed the accuracy of the bone-implant fitting by measuring over and under-hang, namely, the distance between the contours of the bone and the implant. Several zones were ignored: (1) the posterior part of the tibial implant, as the implant shape voluntarily does not follow the resection one; (2) the femoral intercondylar notch, as it is not relevant in the residual pain outcome of the surgery; and (3) the femoral anterior proximal part, as the implant does not try to reproduce anatomy in this zone and, in some cases, the outcrop of the anterior resection plane generated a two bumps profile (see Figure 5).
Figure 5: Flattened contours of the bone resection (in red) and the implant (in blue) for over and under-hang computation. Grey zones are ignored in the computation.
@@ 3.1 Segmentation of partial CT image
The ankle, knee and hip CT image were automatically segmented with a mean DC of $98.5\pm 0.8\%$, mean RMSE of $0.4\pm 0.2mm$ and mean HD of $3.0\pm 1.8mm$. The segmentation accuracy for each bone is detailed in Table 1. The computation time to fully segment all three joints of a patient’s leg was $118\pm 5$ seconds. on CPU (Intel® Coretm i7-7820HQ, 2.90GHz), depending on volume depth.
Bone Extremity DC RMSE HD
Femur Hip $98.6\pm 0.9\%$ $[93.7-99.3]$ $0.4\pm 0.2mm$ $[0.3-1.8]$ $3.0\pm 1.7mm$ $[1.3-10.9]$
Femur Knee $98.6\pm 0.9\%$ $[95.5-99.4]$ $0.4\pm 0.2mm$ $[0.3-1.2]$ $2.8\pm 1.9mm$ $[0.8-9.3]$
Tibia Knee $98.7\pm 0.5\%$ $[96.8-99.4]$ $0.4\pm 0.1mm$ $[0.3-0.8]$ $3.0\pm 1.3mm$ $[1.1-6.9]$
Tibia Ankle $98.2\pm 0.9\%$ $[94.2-99.1]$ $0.4\pm 0.3mm$ $[0.3-1.7]$ $3.1\pm 2.3mm$ $[1.0-9.5]$
Table 1: Segmentation accuracy in terms of Dice coefficient (DC), root mean square error (RMSE) and Hausdorff distance (HD). Results are reported as mean $\pm$ std.dev. [min - max].
@@ 3.2.1 SSM quality
The evaluation metrics of the femur and tibia SSMs are detailed in Table 2, with respect to the number of principal components (modes) kept by the model.
The developed SSMs were able to reconstruct the full femur and tibia shapes with an RMSE of $1.2\pm 0.4mm$ and HD of $5.1\pm 1.9mm$. The tibial SSM fitted the tibia proximal and distal extremities with an RMSE $1.0\pm 0.3mm$ and HD of $4.7\pm 1.7mm$. The femoral SSM fitted the femur proximal and distal extremities with an RMSE of $1.3\pm 0.4mm$ and HD of $5.6\pm 2.0mm$. The reconstruction of the femur and tibia lasted $65\pm 25$ seconds.
Bone Number of modes Compactness Generality Specificity
Tibia 1 88.7% $2.11\pm 0.83mm$ $1.39\pm 0.21mm$
Tibia 3 95.9% $1.72\pm 0.82mm$ $1.42\pm 0.32mm$
Tibia 11 99.0% $1.37\pm 0.77mm$ $1.84\pm 0.70mm$
Femur 1 84.4% $2.30\pm 0.78mm$ $1.74\pm 0.26mm$
Femur 5 95.8% $1.76\pm 0.54mm$ $2.39\pm 0.60mm$
Femur 15 99.1% $1.41\pm 0.68mm$ $2.55\pm 0.49mm$
Table 2: Evaluation metrics of the tibia and femur SSMs.
@@ 3.2.2 Anatomical landmarks determination
The 11 femoral anatomical landmarks have been determined with an accuracy of $2.83\pm 2.15mm$. The accuracy per landmark is detailed in Table 3.
Landmark label mean $\pm$ std [min - max]
AP Sizing Point $3.26\pm 2.05$ $[0.50-9.64]$
Lateral Anterior $2.38\pm 1.58$ $[0.38-7.67]$
Lateral Distal $5.71\pm 3.58$ $[1.29-18.48]$
Lateral Epicondyle $2.47\pm 1.61$ $[0.44-7.55]$
Lateral Posterior $2.67\pm 1.49$ $[0.48-5.78]$
Medial Anterior $2.07\pm 1.89$ $[0.40-11.06]$
Medial Distal $3.21\pm 2.34$ $[0.09-11.05]$
Medial Epicondyle $3.42\pm 1.87$ $[0.46-7.12]$
Medial Posterior $1.91\pm 0.87$ $[0.54-4.23]$
Top Groove $2.00\pm 1.19$ $[0.33-5.49]$
Top Notch $1.99\pm 0.90$ $[0.26-3.61]$
Table 3: Accuracy of the morphological parameters automated determination. The presented errors are the distances between the manual ground truth and the automated computation of the 11 femoral anatomical landmarks.
@@ 3.3 Implant design
Of the 40 testing cases, the custom implant design succeeded in 34 cases (85%) and failed in 6 cases (2 femoral and 1 tibial implants could not be generated, 3 femoral implants and 1 PE insert were invalid). The computation of the custom design for both femoral and tibial implants lasted $23\pm 2$ seconds. The implants fitted the bones with an RMSE of $0.6\pm 0.2mm$ (over and under hang). More details on the implant-to-bone fit can be found in Table 4.
Bone RMSE HD
Femur $0.63\pm 0.26mm$ $2.75\pm 6.19mm$
Tibia $0.65\pm 0.20mm$ $1.22\pm 0.54mm$
Table 4: Bone coverage (over and under hang) for the custom design implants.
@@ 4 Discussion
The proposed method automatically designs patient-specific TKA implants in about 5 minutes, mimicking the joint anatomy and respecting the limb kinematics, to ensure the proper motion and stability of the knee.
@@
The mean segmentation accuracy was $0.4\pm 0.2mm$. We observed high HD on the ankle often resulting from inaccurate detection of the medial malleolus extremity. However, the absence of this portion of the tibial bone does not significantly affect the subsequent planning process. Indeed, the ankle’s primary role is to help determine the mechanical axis of the tibia, which runs from the center of the tibia’s knee to the distal tibial plafond articular surface. This axis orientation remains relatively consistent even if the statistical shape model fits a slightly shorter length for the ankle portion.
Likewise, we observed high HD for segmentation accuracy in the hip region when the great trochanter’s extremity was inaccurately detected. However, these inaccuracies does not significantly impact the positioning of the mechanical axis of the femur after fitting the SSM. (The femoral mechanical axis runs from the center of the femoral head to the knee center located on the distal portion of the femur.)
The segmentation accuracy of our method is in line with similar studies in the literature. For instance, other U-Net-derived methods, for knee segmentation on magnetic resonance images (MRI) or on CT images, reach a mean DC around 98% [44, 45, 46] with either bigger or smaller training and testing dataset.
@@
The proposed method automatically determines anatomical landmarks with a mean accuracy of $2.83\pm 2.15mm$. Such errors are similar to the intra and inter operator variability existing when manually picking the landmarks [39], indicating that the proposed workflow is as accurate as manual acquisitions. The highest errors were obtained for the lateral distal point. This landmark was found to be the hardest to identify as it is situated in a flat area. In the proposed workflow, the lateral distal point is used to determine the distal resection plane only. However, the flat area around the lateral distal point is parallel to the distal resection plane. Therefore, errors in this point detection have negligible impact on the plane orientation and the resulting implant design.
@@
We chose to design our implant following the medial pivot concept which models the femoral condyles with two spheres, thus approximating the sagittal curvature with only one radius for each condyle. The implants design plays an essential role in joint kinematics, and consequently in restoring stable and natural knee movements. For instance, the femoral condyles curvature is a key factor as it impacts the tibiofemoral internal-external rotation and anterior-posterior translation [54]. In this regard the J-curve design has been introduced to efficiently reproduce the femoral condyles sagittal curvature, with different radii of curvature. While the J-curve approach is by essence more anatomical than the medial pivot one, the latter has shown a greater ability to reproduce the natural knee biomechanics [55]. Even though the medial pivot design is still rarely used in practice [56] it is gaining more and more interest, with an increasing number of research articles being published. Recent reviews show that medial pivot TKA is a reliable long-term treatment option for individuals with end-stage osteoarthritis, offering exceptional survivorship, low complication rates, and notable enhancements in clinical and functional outcomes [57, 58].
@@
To our knowledge, only one article in the literature also proposed an automatic workflow to generate personalised implants for TKA from CT images [60]. In their paper, Burge et al. proposed a similar pipeline to ours, using machine learning to segment both the femur and the tibia, and statistical shape modelling to reconstruct the bones 3D models. They obtained custom implants models in less than 10 minutes, by restraining the SSM fitting to very smooth deformations and without determining any anatomical landmark. However, the custom implants they proposed may not be viable in the clinics. Indeed, their implants are directly carved out of the bones 3D models, the implant outer layer corresponding to the surface of the trimmed bone. Such implants replicate the anatomy of the bones – whatever the deformities present – but do not consider any kinematics aspect of the joint, such as the lower limb alignment, the knee flexion axis or the condyles congruence. The personalised implants we propose, are designed regarding more than 70 parameters describing the lower limb anatomy and kinematics, and hence, are better suited for a clinical usage.
@@
It is important to note that the dataset used in our study contains limited instances of severe osteoarthritis, which may not be representative of the patients undergoing total knee arthroplasty. Therefore, the training dataset should be increased to include CT images of such patients to better establish the workflow applicability in the context of surgical planning.
A total of 77 morphological parameters are automatically determined on the patient’s femur and tibia. The accuracy of this morphological analysis was conducted on 11 femoral landmarks only. Although the accuracy of the primary parameters is only necessary (secondary parameters are derived from the primary ones), further work is needed to validate the accuracy of the tibial landmarks too. Nevertheless, as the same methods are used to identify femoral and tibial landmarks, the accuracy for both should be similar.
The proposed workflow does not include the production of the custom designed implants. Such step could be performed in a time-effective manner using 3D printing techniques [61]. However, whatever the techniques used, manufacturing the implants usually involves a third-party supplier, which may add delays between the pre-operative planning and the surgery. Nevertheless, the manual design step is the most time-consuming step that could be addressed with the proposed workflow.
Such personalised prostheses require a higher accuracy than off-the-shelf prostheses when positioning the implants. Indeed, the implants are designed together with the pre-operative planning, ensuring the perfect fit of the implants to the operated bones. During preliminary experiments on cadavers, we observed that inaccuracies on the cutting planes’ angles would affect the quality of implant-to-bone fitting. Patient-specific cutting guides as well as computer assisted solutions (robotic arm or navigated instruments) could help improve the surgical accuracy and ensure the right positioning of the personalised implants.
Eventually, a mechanical evaluation is missing to ensure the stability and survivorship of the personalised implants and to validate the design of the polyethylene tibial insert. Work is currently ongoing to evaluate the implants resistance to fatigue and meet the regulatory standards for future commercialisation.
@@
The workflow can be easily adapted to other joint arthroplasties, after identifying the relevant anatomical landmarks to determine the mechanical and anatomical axes of the concerned joint.

### v3

@@ 2.5.1 Femoral implant
The design of the femoral implant starts with the posterior part and in first place the medial pivot sphere (Fig 3a). For each condyle, the radius of the medial pivot sphere is computed from the antero-posterior size [43]. The posterior surfaces of both medial and lateral condyles are initially shaped based on this sphere, after which they are refined to tightly align with the contours of the resected bone. The posterior most proximal part is then cut along a circle arc to obtain the roll back shape that will allow deep flexion. Secondly, the anterior surface is designed to closely adhere to the patient’s premorbid anatomy, aiming to facilitate patellar tracking. The specific characteristics of each patient’s lateral and medial facet height and sulcus angle are individually taken into consideration to prevent the risk of patellar instability [44, 45].
@@ 2.5.2 Tibial implant
The shape of the tibial baseplate is crafted to conform to the contours of the resected bone, ensuring optimal contact with the cortical bone to enhance implant stability and durability (Fig 3b). The location of the keel is determined by taking into account the morphology of the diaphysis and the tibial plateau. The length of the stem, as well as the dimensions of the two vanes, are also functions of the patient’s specific characteristics.
@@
Fig 3: Design steps for the personalised prosthesis. a) Femoral implant, from left to right: defining the medial pivot sphere; defining the condyles’ curvature; defining the anterior surface; cutting the intercondylar notch; adding pins and chamfers. b) Tibial implant, from left to right: tibial baseplate contouring; adding locking mechanism for the polyethylene insert; adding the keel; adding chamfers.
@@ 2.5.3 Tibial insert
The polyethylene tibial insert is designed regarding the femoral and tibial implants. The basic shape of the insert is built from the embase shape of the tibial component. The footprint of the medial pivot sphere is first excavated into the basic shape. Then, the lateral part is designed almost flat, to allow free antero-posterior translation of the femoral component as suggested by the medial-pivot concept. Finally the most anterior part is cut to avoid conflict with the patella in deep flexion.
@@ 2.5.4 Design evaluation
To evaluate the quality of the designed custom implants, the accuracy of the bone-implant fitting is computed by measuring over and under-hang, namely, the distance between the contours of the bone and the implant. Several zones were ignored: (1) the posterior part of the tibial implant, as the implant shape voluntarily does not follow the resection one; (2) the femoral intercondylar notch, as it is not relevant in the residual pain outcome of the surgery; and (3) the femoral anterior proximal part, as the implant does not try to reproduce the anatomy in this zone (see Fig 4).
Fig 4: Flattened contours used for over and under-hang computation. The red and blue lines represent bone resection and implant contours respectively. Grey zones are ignored in the computation.
@@ 3.1 Segmentation of partial CT image
The ankle, knee and hip CT image were automatically segmented with a mean DC of $98.5\pm 0.8\%$, mean RMSE of $0.4\pm 0.3mm$ and mean HD of $3.0\pm 2.1mm$. The segmentation accuracy for each bone is detailed in Table 1. The computation time to fully segment all three joints of a patient’s leg was $378\pm 221$ seconds on a CPU (Intel® Coretm i7-7820HQ, 2.90GHz), depending on the volume depth.
{adjustwidth}
-2.25in0in Bone Extremity DC (%) RMSE (mm) HD (mm) Femur Proximal $98.6\pm 0.9$ $[93.7-99.3]$ $0.4\pm 0.2$ $[0.3-1.8]$ $3.0\pm 1.7$ $[1.3-10.9]$ Distal $98.6\pm 0.9$ $[95.5-99.4]$ $0.4\pm 0.4$ $[0.2-2.4]$ $3.1\pm 2.8$ $[0.9-15.1]$ Tibia Proximal $98.7\pm 0.5$ $[96.8-99.4]$ $0.3\pm 0.1$ $[0.0-0.5]$ $2.8\pm 1.3$ $[0.0-5.8]$ Distal $98.2\pm 0.9$ $[94.2-99.1]$ $0.4\pm 0.3$ $[0.0-1.7]$ $3.0\pm 2.4$ $[0.0-9.5]$
Dice coefficient (DC), root mean square error (RMSE) and Hausdorff distance (HD) are reported as mean $\pm$ std.dev. [min - max].
Table 1: Segmentation accuracy.
@@ 3.2.1 SSM quality
The evaluation metrics of the femur and tibia SSMs are detailed in Table 2, with respect to the number of principal components (modes) kept by the model. The SSMs are later reduced to their first N modes of deformation representing at least 99% of variability to simplify and speed up the SSMs fitting (i.e. 15 and 11 modes for the femur and tibia SSMs respectively).
Table 2: Tibia and femur SSMs metrics, regarding the number of modes (N).
Bone N Compactness (%) Generality (mm) Specificity (mm)
Tibia 1 88.7 $2.11\pm 0.83$ $1.39\pm 0.21$
3 95.9 $1.72\pm 0.82$ $1.42\pm 0.32$
11 99.0 $1.37\pm 0.77$ $1.84\pm 0.70$
89 100 $1.26\pm 0.39$ $2.90\pm 0.90$
Femur 1 84.4 $2.30\pm 0.78$ $1.74\pm 0.26$
5 95.8 $1.76\pm 0.54$ $2.39\pm 0.60$
15 99.1 $1.41\pm 0.68$ $2.55\pm 0.49$
89 100 $1.31\pm 0.79$ $1.84\pm 0.41$
Generality and specificity are reported as mean $\pm$ std.dev.
3.2.2 Bone reconstruction
The full femur and tibia bones were automatically reconstructed from partial data with a mean RMSE of $1.0\pm 0.3mm$ and mean HD of $4.1\pm 1.3mm$. The registration accuracy for each bone is detailed in Table 3. The reconstruction of the femur and tibia lasted $329\pm 55$ seconds.
Table 3: Full bone reconstruction accuracy.
Bone Extremity RMSE (mm) HD (mm)
Femur Proximal $1.2\pm 0.2$ $[0.9-1.5]$ $4.5\pm 0.9$ $[3.2-6.6]$
Distal $1.0\pm 0.3$ $[0.7-2.1]$ $4.1\pm 1.3$ $[2.7-8.8]$
Tibia Proximal $1.0\pm 0.3$ $[0.7-2.3]$ $4.1\pm 1.2$ $[2.8-9.1]$
Distal $0.8\pm 0.3$ $[0.5-2.1]$ $3.6\pm 1.7$ $[1.9-10.5]$
RMSE and HD are reported as mean $\pm$ std.dev. [min - max].
@@ 3.2.3 Anatomical landmarks determination
The anatomical landmarks have been determined with a mean accuracy of $2.2\pm 1.5mm$ ($2.5\pm 1.6mm$ and $1.5\pm 1.2mm$ for the femoral and tibial landmarks respectively). The accuracy per landmark is detailed in Table 4.
{adjustwidth}
-2.25in0in Femoral Landmarks mean $\pm$ std [min - max] AP Sizing Point $3.0\pm 1.7$ $[0.3-8.2]$ Lateral Anterior $1.9\pm 1.2$ $[0.2-5.3]$ Lateral Distal $4.0\pm 1.8$ $[0.8-7.9]$ Lateral Epicondyle $2.2\pm 1.2$ $[0.5-6.2]$ Lateral Posterior $2.8\pm 1.4$ $[0.3-6.3]$ Medial Anterior $1.3\pm 0.8$ $[0.3-4.0]$ Medial Distal $3.1\pm 1.8$ $[0.2-9.3]$ Medial Epicondyle $3.0\pm 1.8$ $[0.2-7.6]$ Medial Posterior $1.8\pm 0.8$ $[0.4-3.4]$ Top Groove $2.2\pm 1.2$ $[0.5-4.8]$ Top Notch $1.9\pm 1.0$ $[0.4-4.7]$ Tibial Landmarks mean $\pm$ std [min - max] Ankle Joint Center $2.5\pm 1.0$ $[0.9-5.1]$ Knee Center $2.2\pm 0.9$ $[0.7-4.1]$ Lateral Wear Point $0.7\pm 0.6$ $[0.0-2.5]$ Medial Wear Point $0.7\pm 0.6$ $[0.0-2.5]$ Condyle Centers Joint Line $2.0\pm 1.6$ $[0.0-6.2]$ Errors in mm, except for the tibial condyle centers line angular error in .
Table 4: Accuracy of the morphological parameters automatic determination.
@@ 3.3 Implant design
Designing the custom implants succeeded in 89% of cases and failed in 4 cases (2 femoral implants, 1 tibial implant and 1 PE insert were invalid). The computation of the custom design for both femoral and tibial implants lasted $270\pm 80$ seconds. The implants fitted the bones with an RMSE of $0.9\pm 0.5mm$ and HD of $3.1\pm 2.1mm$ (over and under hang). More details on the implant-to-bone fitting can be found in Table 5. The highest hausdorff values corresponded to under hang at the posterior lateral tibial plateau. Such coverage was obtained due to a retreat of the plateau in the design to avoid muscle impingement.
The custom implants for two patients are illustrated in Fig 5.
Table 5: Bone coverage for custom design implants (over and under hang).
Bone RMSE (mm) HD (mm)
Femur $0.4\pm 0.1$ $[0.3-0.6]$ $1.2\pm 0.4$ $[0.7-2.8]$
Tibia $1.4\pm 0.3$ $[0.9-1.9]$ $4.9\pm 1.3$ $[2.5-7.1]$
RMSE and HD are expressed as mean $\pm$ std.dev. [min - max].
Anterior view   Posterior view   Sagittal view
Refer to captionRefer to captionRefer to caption
Fig 5: Example of personalised implants designed for two patients for their right (top row) and left (bottom row) knees.
@@ 4 Discussion
The proposed method automatically designs patient-specific TKA implants in about 15 minutes, mimicking the joint anatomy and respecting the limb kinematics, by using the medial pivot prosthetic concept, to ensure the proper motion and stability of the knee.
@@
The mean segmentation accuracy was $0.4\pm 0.3mm$. We observed high HD on the ankle often resulting from inaccurate detection of the medial malleolus extremity. However, the absence of this portion of the tibial bone does not significantly affect the subsequent planning process. Indeed, the ankle’s primary role is to help determine the mechanical axis of the tibia, which runs from the knee center located on the proximal tibia to the distal tibial plafond articular surface. This axis orientation remains relatively consistent even if the SSM fits a slightly shorter length for the ankle portion.
Likewise, we observed high HD for segmentation accuracy in the hip region when the great trochanter’s extremity was inaccurately detected. However, these inaccuracies does not significantly impact the positioning of the mechanical axis of the femur after fitting the SSM. (The femoral mechanical axis runs from the center of the femoral head to the knee center located on the distal femur.)
The segmentation accuracy of our method is in line with similar studies in the literature. For instance, other U-Net derived methods, for knee segmentation on magnetic resonance images (MRI) or on CT images, reach a mean DC around 98% with either bigger or smaller training and testing dataset [46, 47, 48].
@@
The proposed method automatically determines anatomical landmarks with a mean accuracy of $2.2\pm 1.5mm$. Such errors are similar to the intra and inter operator variability existing when manually picking the landmarks [41], indicating that the proposed workflow is as accurate as manual acquisitions. The highest errors were obtained for the lateral distal point (most distal point on the femoral lateral condyle). This landmark was found to be the hardest to identify as it is situated in a flat area. In the proposed workflow, the lateral distal point is used to determine the femoral distal resection plane only. However, the flat area around the lateral distal point is parallel to the this resection plane. Therefore, errors in this point detection have negligible impact on the resection plane orientation and the resulting implant design.
@@
We chose to design our implant following the medial pivot concept which models the femoral condyles with two spheres, thus approximating the sagittal curvature with only one radius for each condyle. The implants design plays an essential role in joint kinematics, and consequently in restoring stable and natural knee movements. For instance, the femoral condyles curvature is a key factor as it impacts the tibiofemoral internal-external rotation and antero-posterior translation [56]. In this regard the J-curve design has been introduced to efficiently reproduce the femoral condyles sagittal curvature, with different radii of curvature. While the J-curve approach is by essence more anatomical than the medial pivot one, the latter has shown a greater ability to reproduce the natural knee biomechanics [57]. Even though the medial pivot design is still rarely used in practice [58] it is gaining more and more interest, with an increasing number of research articles being published. Recent reviews show that medial pivot TKA is a reliable long-term treatment option for individuals with end-stage osteoarthritis, offering exceptional survivorship, low complication rates, and notable enhancements in clinical and functional outcomes [59, 60].
@@
To our knowledge, only one article in the literature also proposed a complete automatic workflow to generate personalised implants for TKA from CT images. Burge et al. [34] recently proposed a similar pipeline to ours, using machine learning to segment both the femur and the tibia, and statistical shape modelling to reconstruct the bones 3D models. They obtained custom implants models in less than 10 minutes, by restraining the SSM fitting to very smooth deformations and without determining any anatomical landmark. However, the custom implants they proposed may not be viable in the clinics. Indeed, their implants are directly carved out of the bones 3D models, the implant outer layer corresponding to the surface of the trimmed bone. Such implants replicate the anatomy of the bones – whatever the deformities present – but do not consider any kinematics aspect of the joint, such as the lower limb alignment, the knee flexion axis or the condyles congruence. The personalised implants we propose, are designed regarding more than 70 parameters describing the lower limb anatomy and kinematics, and hence, are better suited for a clinical usage.
Semi automated procedures have been proposed in the literature to shorten segmentation, pre-operative planning or implant design [33, 62, 63], but the required manual interventions remain tedious and time-consuming which limits their usability in clinical routine. Today, customisation solutions available on the market are still very little used. For instance, 10 000 custom implants from Symbios Orthopédie SA were implanted between 2018 and 2022, which represent less than 3% of the TKAs in France on the same period (more than 300 000)[13, 64]. Consequently, automating and simplifying all the processes from segmentation to implant manufacturing is needed to perform personalised TKA routinely. The proposed approach for automating CT images segmentation, morphological analysis, and implant design, not only reduces the time needed for preparing custom TKA, but also enables surgeons to control the entire planning process, from start to finish, and to fine-tune it within a minute if necessary. Such automation simplifies the planning customisation, as it reduces the need to go back and forth with the implant designers and manufacturers to test different configurations, modifying one parameter at a time.
@@
It is important to note that the dataset used in our study contains limited instances of severe osteoarthritis, which may not be representative of all patients undergoing TKA. The inclusion criteria were designed to be not too restrictive in order to recruit a large variety of patients. However, it was collected from only one center which may be representative of a local population rather than the global population. Therefore, the training and testing dataset should be increased to include CT images of more diverse patients to better establish the workflow applicability to a broader population in the context of surgical planning.
A total of 77 morphological parameters are automatically determined on the patient’s femur and tibia. The accuracy of this morphological analysis was conducted on 17 femoral and tibial primary landmarks only. The remaining primary landmarks, such as the femoral head center, were not considered here as they cannot be picked manually on 3D printed or virtual meshes. The accuracy of secondary parameters was not analysed here, as secondary parameters are derived from primary ones and are directly link to the primary landmarks accuracy. Moreover, secondary parameters’ ground truth may vary according to the surgeon preferences and the surgical plan adopted (e.g. resection plans orientation and depth).
The proposed workflow does not include the production of the custom designed implants. Such step could be performed in a time-effective manner using 3D printing techniques [65]. However, whatever the technique used, manufacturing the implants usually involves a third-party supplier, which may add delays between the pre-operative planning and the surgery. Nevertheless, designing custom implants manually remains a time-consuming task that could be swiftly addressed with the proposed workflow.
Such personalised prostheses require a higher accuracy than off-the-shelf prostheses when positioning the implants. Indeed, the implants are designed together with the pre-operative planning, ensuring the perfect fit of the implants to the operated bones. During preliminary experiments on cadavers, we observed that inaccuracies on the cutting planes’ positioning would affect the quality of implant-to-bone fitting. Patient-specific cutting guides as well as computer assisted solutions (robotic arm or navigated instruments) could help improve the surgical accuracy and ensure the right positioning of the personalised implants.
Finally, a mechanical evaluation is missing to ensure the stability and survivorship of the personalised implants and to validate the design of the polyethylene tibial insert. Work is currently ongoing to evaluate the implants resistance to fatigue and meet the regulatory standards for future commercialisation.
@@
The proposed workflow can be easily adapted to other joint arthroplasties, once the relevant anatomical landmarks are identified.
Finally, a cost analysis will be carried out in the coming years to validate the feasibility of the proposed solution and subsequently integrate it into clinical practice. To this day, current solutions for producing custom implants require engineering work to process the CT images, model the patient’s anatomy and design the implants. Such work represent a significant cost in the manufacture of implants, which may vary from a company to the other, and a country to the other. The proposed solution automates these processing, modelling and designing steps, considerably reducing the time and engineering work required to carry them out. As a result, the associated costs will also fall significantly.